# In search of the electron-phonon contribution to total energy

### v1

@@ III Mass-scaling analysis
In the original BO work [Born1927], it is shown how to formulate a perturbation expansion around a minimum (global or local) of a chosen BO energy hypersurface with index $j$, taken as reference configuration, when the nuclei masses are much larger than the electronic masses. The small parameter of this expansion is here denoted $\lambda$. Taking $M_{0}$ as a characteristic nuclear mass, expressed in units of electronic mass, one defines relative nuclear masses
@@
Quantities that depend on $\lambda$ (the wavefunctions, Hamiltonian, total energy, and Green’s function) can be expanded in a Taylor series in $\lambda$ as follows, generically:
@@
Figure 2: Same coordinate diagram as Fig. 1 where we focus on the specific $j=0$ BO hypersurface (black curved line). This BO hypersurface presents here two minima denoted $\{\mathbf{R}_{j=0}^{0}\}$ and $\{\mathbf{R}_{j=0}^{0^{\prime}}\}$. The exact total energy $\mathcal{E}_{i}$ (black horizontal line) of some level $i$ is obtained from a series expansion considering one or the other reference configuration as starting point, in the neighborhood of which the BO hypersurface is expanded to quadratic order (red-dashed parabola), defining an harmonic oscillator Hamiltonian. An eigenstate of this harmonic oscillator Hamiltonian is labeled with the index $\jmath$ (dotless-$j$). The corresponding eigenenergy is added to the reference BO energy giving the red dashed horizontal line. The series expansion (red dash) maps the corresponding black energy level if the series converges.
@@
$\hat{\mathfrak{H}}_{j}^{[0]}\{{\mathbf{R}}\}\equiv\sum_{\kappa\alpha}\frac{\hat{P}^{2}_{\kappa\alpha}}{2m_{\kappa}}+E^{\rm BO\{\mathbf{2}\}}_{j}\cdot\Delta\mathbf{R}_{j}^{2},$ (45)
with associated nuclear wavefunctions, $\vartheta_{\jmath j}$, and eigenenergies $\mathfrak{E}_{\jmath j}^{[0]}$,
$\hat{\mathfrak{H}}_{j}^{[0]}\{{\mathbf{R}}\}\vartheta_{\jmath j}\{{\mathbf{R}}\}=\mathfrak{E}_{\jmath j}^{[0]}\vartheta_{\jmath j}\{{\mathbf{R}}\},$ (46)
where the subscript $\jmath$ (dotless $j$) characterizes the nuclear state, considering the $j$ BO hypersurface energy and curvature in the chosen reference configuration. The expression of the energy in terms of the nuclear state and the expansion around a (local or global) minima is depicted in Fig. 2 for the specific $j=0$ BO hypersurface, for which one has to choose among two possible reference configurations. Importantly, the nuclear state index $\jmath$ starts at zero around each minimum $\{\mathbf{R}_{j}^{0}\}$ of the $j$ hypersurface. Both $\hat{\mathfrak{H}}_{j}^{[0]}\{{\mathbf{R}}\}$ and $\mathfrak{E}_{\jmath j}^{[0]}$ are denoted with a gothic calligraphy to indicate that they are related only to nuclear dynamics. The reason for the superscript $[0]$ of the Hamiltonian will be explained later, see Eq. (77). Because $\hat{\mathfrak{H}}_{j}^{[0]}$ is even with respect to a sign change of $\Delta{\mathbf{R}}_{j}$, its eigenfunctions are even or odd under such sign change.
In Appendix B, we show that the Schrödinger equation, Eq. (38), is satisfied up to the second order in the displacements by the following wavefunction,
@@
that has the form of Eq. (22b) and where we have associated the index $i=(\jmath,j)$ to a specific pair $\jmath$ and $j$. This association is exact when the harmonic oscillator series converges, see Fig. 2. The nuclear displacements are scaled by the inverse of $\lambda$ in the first factor, describing the nuclear dynamics, while they are not scaled in the second factor, describing the electronic BO wavefunctions. In Appendix B, it is also shown that $\mathcal{E}^{(2)}_{i}=\mathfrak{E}_{\jmath j}^{[0]}$, so that
$\mathcal{E}_{i\lambda}=E^{\rm BO}_{j|\mathbf{0}}+\lambda^{2}\mathfrak{E}_{\jmath j}^{[0]}+\mathcal{O}(\lambda^{3}),$ (48)
where the first term is independent of the characteristic mass and the second term scales as $M_{0}^{-1/2}$. In Eq. (48), the second-order energy is the correction with respect to a zero-order energy that depends on the index $j$.
Born and Oppenheimer [Born1927] have analyzed anharmonic corrections, expanding $E_{j}^{\rm BO}\{{\mathbf{R}}\}$ beyond the second order in the nuclear displacements, up to the fourth order in $\lambda$. We proceed with the analysis of the mass scaling of other corrections beyond those examined by them, up to the sixth order in $\lambda$. In order to do this, inspired by Eq. (47) we define scaled nuclear displacements with respect to the reference configuration, and express the nuclear dynamics in terms of these. We define
@@
Importantly, the derivatives of the electronic BO wavefunctions present in $A_{\kappa\alpha j}\{{\mathbf{R}}\}$, Eq. (11), in $B_{\perp\kappa\alpha j}\{{\mathbf{R}}\}$, Eq. (17), as well as in $\hat{\cal{T}}^{\textrm{N}}_{j}\{{\mathbf{R}}\}$, Eq. (27), remain derivatives in terms of the nuclear configurations. This is due to the expression of the second factor in Eq. (47), which is still expressed in nuclear configurations, not in scaled displacements. However, these functions are evaluated in the configuration expressed equivalently by $\{{\mathbf{R}}\}$ or by $\lambda\boldsymbol{\rho}]$. So, one has
@@
We now examine the mass-scaling behavior of the different approximations presented in Sec. II. Similarly to Eqs. (22a)-(22d), we express the total wavefunction using scaled displacements $\boldsymbol{\rho}_{j}]$ instead of real space coordinates:
@@
In the scalar approximation, one sets $\lambda A_{\kappa\alpha j}[\lambda\boldsymbol{\rho}_{j}]$ and $E^{\rm Born}_{j}[\lambda\boldsymbol{\rho}_{j}]$ to zero, while $\lambda A_{\kappa\alpha j}[\lambda\boldsymbol{\rho}_{j}]$ is retained in the “hyb1” approximation, and $E^{\rm Born}_{j}[\lambda\boldsymbol{\rho}_{j}]$ is retained in the “hyb2” approximation. The exact equation for the energy $\mathcal{E}_{i\lambda}$ of the full Hamiltonian, Eq. (57), is now expressed in term of the self-energy, Eq. (30):
@@
In Eq. (61), the reference BO energy $E^{\rm BO}_{j|\textbf{0}}$ can be subtracted from both $E^{\rm opt}_{j\lambda}[\boldsymbol{\rho}_{j}]$ and $\mathcal{E}_{i\lambda}$, leaving quantities that are at least second-order in $\lambda$. Such second-order component can be used as a zeroth-order Hamiltonian to find the remaining expansion for $\mathcal{E}_{i\lambda}$ and to find the wavefunction $\chi_{\jmath j\lambda}[\boldsymbol{\rho}_{j}]$. Explicitly, Eq. (61) becomes
$\Big(\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]+\hat{\mathfrak{H}}^{[1]}_{j\lambda}[\boldsymbol{\rho}_{j}]\Big)\chi_{\jmath j\lambda}[\boldsymbol{\rho}_{j}]=\mathfrak{E}_{\jmath j\lambda}\chi_{\jmath j\lambda}[\boldsymbol{\rho}_{j}],$ (66)
@@
$\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]\equiv\sum_{\kappa\alpha}\frac{\hat{\pi}^{2}_{\kappa\alpha}}{2m_{\kappa}}+E^{\rm BO\{\mathbf{2}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{2},$ (67)
and corrections to this Hamiltonian
$\hat{\mathfrak{H}}_{j\lambda}^{[1]}[\boldsymbol{\rho}_{j}]\equiv$ $\lambda\hat{T}_{j}^{\textrm{vec(1)}}[\lambda\boldsymbol{\rho}_{j}]+\lambda^{2}\hat{T}_{j}^{\textrm{vec(2)}}[\lambda\boldsymbol{\rho}_{j}]$
@@
The eigenvalues in Eq. (66) are related to those in Eq. (61) by
Without the correction $\hat{\mathfrak{H}}_{j\lambda}^{[1]}$, Eq. (66) reduces to the $\lambda$-independent equation
$\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]\chi_{\jmath j}^{[0]}[\boldsymbol{\rho}_{j}]=\mathfrak{E}_{\jmath j}^{[0]}\chi_{\jmath j}^{[0]}[\boldsymbol{\rho}_{j}],$ (73)
@@
$\chi_{\jmath j}^{[0]}[\boldsymbol{\rho}_{j}]=\vartheta_{\jmath j}[\boldsymbol{\rho}_{j}].$ (74)
Contributions with different powers of $\lambda$ in $\hat{\mathfrak{H}}_{j\lambda}^{[1]}[\boldsymbol{\rho}_{j}]$ can be ordered to understand the effect of ${\cal{O}}(\lambda)$ and ${\cal{O}}(\lambda^{2})$ contributions on $\mathfrak{E}_{\jmath j\lambda}$. Moreover, it is possible to fix the gauge of the BO wavefunction at the reference configuration, as well as its first-order derivative with respect to atomic displacements, such that $A_{\kappa\alpha j}[\boldsymbol{0}]=0$. Hence $\hat{T}_{j}^{\textrm{vec(1)}}$ is ${\cal{O}}(\lambda)$ and $\hat{T}_{j}^{\textrm{vec(2)}}$ is ${\cal{O}}(\lambda^{2})$. Equation (III) becomes
$\hat{\mathfrak{H}}_{j\lambda}^{[1]}[\boldsymbol{\rho}_{j}]\equiv\lambda E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}$
@@
We treat Eq. (66) with perturbation theory and write a new expansion in terms of the following $\Lambda$-dependent Hamiltonian
$\hat{\mathfrak{H}}_{j\lambda}([\boldsymbol{\rho}_{j}],\Lambda)=\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]+\Lambda\hat{\mathfrak{H}}^{[1]}_{j\lambda}[\boldsymbol{\rho}_{j}],$ (76)
with the generic perturbation series
$X(\Lambda)\equiv X^{[0]}+\Lambda X^{[1]}+\Lambda^{2}X^{[2]}+\mathcal{O}(\Lambda^{3}),$ (77)
and $\Lambda$ as a small parameter. For the energy, one has
$\mathcal{E}_{i\lambda}-E_{j|\textbf{0}}^{\textrm{BO}}=\lambda^{2}\Big(\mathfrak{E}_{\jmath j}^{[0]}+\Lambda\mathfrak{E}_{\jmath j\lambda}^{[1]}+\Lambda^{2}\mathfrak{E}_{\jmath j\lambda}^{[2]}\\ +\Lambda^{3}\mathfrak{E}_{\jmath j\lambda}^{[3]}+\Lambda^{4}\mathfrak{E}_{\jmath j\lambda}^{[4]}+\mathcal{O}(\Lambda^{5})\Big),$ (78)
where $\mathcal{E}_{i}^{(2)}=\mathfrak{E}_{\jmath j}^{[0]}$ is independent of $\Lambda$.
@@
As previously observed, $\vartheta_{\jmath j}$ is even or odd with respect to a change of sign of $\boldsymbol{\rho}_{j}$ (also changing sign of the associated momentum). Moreover, all terms of $\hat{\mathfrak{H}}^{[1]}_{j\lambda}$ that are even with respect to such a sign change have an even $\lambda$ exponent, while all terms that are odd with respect to such a sign change have an odd $\lambda$ exponent. Hence, the $\lambda$-expansion of the total energy turns out to have only even power terms.
At this stage, let us examine the contributions to $\mathcal{E}_{i\lambda}-E_{j|\textbf{0}}^{\textrm{BO}}$ up to order five in $\lambda$. We already have the zero- and second-order contributions, see Eqs. (46)-(48), while the odd-$\lambda$ terms vanish. So, we should compute $\mathfrak{E}_{\jmath j}^{(2)}$, which amounts to treating the effect of $\hat{\mathfrak{H}}_{j\lambda}^{[1]}$ up to second order in $\Lambda$. As usual in perturbation theory, the Hellmann-Feynman theorem allows one to compute the first-order contribution of $\hat{\mathfrak{H}}_{j\lambda}^{[1]}$ in $\Lambda$ by taking the expectation value of the Hamiltonian correction for the unperturbed wavefunction,
$\mathfrak{E}_{\jmath j\lambda}^{[1]}=\langle\vartheta_{\jmath j}|\hat{\mathfrak{H}}^{[1]}_{j\lambda}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}=\lambda^{2}\mathfrak{E}_{\jmath j}^{[1](2)}+\mathcal{O}(\lambda^{4}),$ (79)
with
$\mathfrak{E}_{\jmath j}^{[1](2)}=$ $\mathcal{E}_{i}^{(4),\text{vec}}+\langle\vartheta_{\jmath j}|E^{\rm BO\{\mathbf{4}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{4}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}$
$+\langle\vartheta_{\jmath j}|E^{\rm Born}_{j}[\bf{0}]|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}.$ (80)
where
$\mathcal{E}_{i}^{(4),\text{vec}}=\langle\vartheta_{\jmath j}|\hat{T}_{j}^{\textrm{vec(1)\{{\bf{1}}\}}}\cdot\boldsymbol{\rho}_{j}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}.$ (81)
The second-order contribution of $\hat{\mathfrak{H}}_{j\lambda}^{[1]}$ in $\Lambda$ can be obtained by a sum-over-state formulation, using a Sternheimer equation, or using a Green’s function approach, like for the case of Eqs. (29)-(31). We will use the latter. We define the Green’s function of the auxiliary harmonic oscillator Hamiltonian $\hat{\mathfrak{H}}_{j}^{[0]}\{{\mathbf{R}}\}$, see Eq. (45), as
$\hat{\mathfrak{G}}_{\perp\jmath j}(z,[\boldsymbol{\rho}_{j}])\equiv\hat{\cal{P}}_{\perp\vartheta_{\jmath j}}\Big(z-\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]\Big)^{-1}\hat{\cal{P}}_{\perp\vartheta_{\jmath j}},$ (82)
with the term between parenthesis being a pseudo-inverse, limited to the subspace perpendicular to $\vartheta_{\jmath j}[\boldsymbol{\rho}_{j}]$. The Green’s function evaluated at $\mathfrak{E}_{\jmath j}^{[0]}$ is labeled with a $[0]$ superscript:
$\hat{\mathfrak{G}}_{\perp\jmath j}^{[0]}[\boldsymbol{\rho}_{j}]\equiv\hat{\mathfrak{G}}_{\perp\jmath j}(\mathfrak{E}_{\jmath j}^{[0]},[\boldsymbol{\rho}_{j}])$ (83)
Then,
$\mathfrak{E}_{\jmath j\lambda}^{[2]}=\langle\vartheta_{\jmath j}|\hat{\mathfrak{H}}^{[1]}_{j\lambda}\hat{\mathfrak{G}}_{\perp\jmath j}^{[0]}[\boldsymbol{\rho}_{j}]\hat{\mathfrak{H}}^{[1]}_{j\lambda}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}.$ (84)
Considering this expression up to second order in $\lambda$ gives
$\mathfrak{E}_{\jmath j\lambda}^{[2]}=$ $\lambda^{2}\mathfrak{E}_{\jmath j}^{[2](2)}+\mathcal{O}(\lambda^{4}),$ (85a)
$\mathfrak{E}_{\jmath j}^{[2](2)}=$ $\langle\vartheta_{\jmath j}|\hat{\mathfrak{H}}^{[1](1)}_{j}\hat{\mathfrak{G}}_{\perp\jmath j}^{[0]}[\boldsymbol{\rho}_{j}]\hat{\mathfrak{H}}^{[1](1)}_{j}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}},$ (85b)
$\hat{\mathfrak{H}}^{[1](1)}_{j}$= $E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}.$ (85c)
With the explicit expressions Eqs. (III) and (85b), the mass-scaling expansion of the total eigenenergies obtained from the different approximations presented in Sec. II can be established. They all share the same zero- and second-order terms, see Eq. (48). The fourth-order terms are the sum of contributions from Eqs. (III) and (85b), with all terms being present for the optimal approximation,
$\mathcal{E}_{i}^{\rm(4)}=\mathfrak{E}_{\jmath j}^{[1](2)}+\mathfrak{E}_{\jmath j}^{[2](2)},$ (86)
which is the same as the expansion of the exact energy $\mathcal{E}_{i}$ up to order 5 included. For the scalar approximation, all vector contributions and the Born contribution are neglected, which gives:
@@
$\mathcal{E}_{i}^{\rm(4),anh}=\langle\vartheta_{\jmath j}|E^{\rm BO\{\mathbf{4}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{4}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}\\ +\langle\vartheta_{\jmath j}|E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}\hat{\mathfrak{G}}_{\perp\jmath j}^{[0]}[\boldsymbol{\rho}_{j}]E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}},$ (87)
that is, only the BO anharmonic terms are kept. For the “hyb1” approximation, the Born contribution in $\mathfrak{E}_{\jmath j}^{[1](2)}$ is neglected, while for the “hyb2”, the vector potential contribution in $\mathfrak{E}_{\jmath j}^{[1](2)}$ is neglected.
@@
Now that we have established the mass scaling of the four adiabatic approximations up to order five, with the definition of an auxiliary harmonic oscillator Hamiltonian, we can carry on our search for the electron-phonon contribution to the total energy.

### v2

@@ III Mass-scaling analysis
In the original BO work [13], it is shown how to formulate a perturbation expansion around a minimum (global or local) of a chosen BO energy hypersurface with index $j$, taken as reference configuration, when the nuclei masses are much larger than the electronic masses. The small parameter of this expansion is here denoted $\lambda$. Taking $M_{0}$ as a characteristic nuclear mass, expressed in units of electronic mass, one defines relative nuclear masses
@@
Quantities that depend on $\lambda$ (wavefunctions, Hamiltonian, total energy, and Green’s function) can be expanded in a Taylor series in $\lambda$ as follows, generically:
@@
Figure 2: Same coordinate diagram as Fig. 1 where we focus on the specific $j=0$ BO hypersurface (black curved line). This BO hypersurface presents here two minima denoted $\{\mathbf{R}_{j=0}^{0}\}$ and $\{\mathbf{R}_{j=0}^{0^{\prime}}\}$. The exact total energy $\mathcal{E}_{i}$ (black horizontal line) of some level $i$ is obtained from a series expansion considering one or the other reference configuration as starting point, in the neighborhood of which the BO hypersurface is expanded to quadratic order (dashed parabola), defining an harmonic oscillator Hamiltonian. An eigenstate of this harmonic oscillator Hamiltonian is labeled with the index $\jmath$ (dotless-$j$). The corresponding eigenenergy is added to the reference BO energy giving the red dashed horizontal line. The series expansion (red dash) maps the corresponding black energy level if the series converges.
@@
$\hat{\mathfrak{H}}_{j}^{\textrm{ho}}\{\Delta{\mathbf{R}}\}\equiv\sum_{\kappa\alpha}\frac{\hat{P}^{2}_{\kappa\alpha}}{2m_{\kappa}}+E^{\rm BO\{\mathbf{2}\}}_{j}\cdot\Delta\mathbf{R}_{j}^{2},$ (46)
with associated nuclear wavefunctions, $\vartheta_{\jmath j}$, and eigenenergies $\mathfrak{E}_{\jmath j}^{\textrm{ho}}$,
$\hat{\mathfrak{H}}_{j}^{\textrm{ho}}\{\Delta{\mathbf{R}}\}\vartheta_{\jmath j}\{\Delta{\mathbf{R}}\}=\mathfrak{E}_{\jmath j}^{\textrm{ho}}\vartheta_{\jmath j}\{\Delta{\mathbf{R}}\},$ (47)
where the subscript $\jmath$ (dotless $j$) characterizes the nuclear state, considering the $j$ BO hypersurface energy and curvature in the chosen reference configuration. The expression of the energy in terms of the nuclear state and the expansion around a (local or global) minima is depicted in Fig. 2 for the specific $j=0$ BO hypersurface, for which one has to choose among two possible reference configurations. Importantly, the nuclear state index $\jmath$ starts at zero around each minimum $\{\mathbf{R}_{j}^{0}\}$ of the $j$ hypersurface. Both $\hat{\mathfrak{H}}_{j}^{\textrm{ho}}\{\Delta{\mathbf{R}}\}$ and $\mathfrak{E}_{\jmath j}^{\textrm{ho}}$ are denoted with a gothic calligraphy to indicate that they are related only to nuclear dynamics. Because $\hat{\mathfrak{H}}_{j}^{\textrm{ho}}$ is even with respect to a sign change of $\Delta{\mathbf{R}}_{j}$, its eigenfunctions are even or odd under such a sign change.
In Sec. LABEL:App:wf_factorizes of the SI, we show that the Schrödinger equation, Eq. (39), is satisfied up to the second order in the displacements by the following wavefunction,
@@
that has the form of Eq. (23b) and where we have associated the index $i=(\jmath,j)$ to a specific pair $\jmath$ and $j$ where the index $\jmath$ is associated with the local minimum $\mathbf{R}_{j}^{0}$} which best describes $i$. This association is possible when the $\lambda$-based expansion starting from the auxiliary harmonic oscillator Hamiltonian converges and is exemplified by Fig. 2. The nuclear displacements are scaled by the inverse of $\lambda$ in the first factor, describing the nuclear dynamics, while they are not scaled in the second factor, describing the electronic BO wavefunctions. In Sec. LABEL:App:wf_factorizes of the SI, it is also shown that $\mathcal{E}^{(2)}_{i}=\mathfrak{E}_{\jmath j}^{\textrm{ho}}$, so that
$\mathcal{E}_{i\lambda}=E^{\rm BO}_{j|\mathbf{0}}+\lambda^{2}\mathfrak{E}_{\jmath j}^{\textrm{ho}}+\mathcal{O}(\lambda^{3}),$ (49)
where the first term is independent of the characteristic mass and the second term scales as $M_{0}^{-1/2}$. In Eq. (49), the second-order energy is the vibrational correction with respect to a zero-order energy that depends on the index $j$.
Born and Oppenheimer [13] have analyzed anharmonic corrections, expanding $E_{j}^{\rm BO}\{{\mathbf{R}}\}$ beyond the second order in the nuclear displacements, up to the fourth order in $\lambda$. We proceed with the analysis of the mass scaling of other corrections beyond those examined by them, up to the sixth order in $\lambda$. In order to do this, inspired by Eq. (48) we define scaled nuclear displacements with respect to the reference configuration, and express the nuclear dynamics in terms of these. Explicitly,
@@
Importantly, the derivatives of the electronic BO wavefunctions present in $A_{\kappa\alpha j}\{{\mathbf{R}}\}$, Eq. (13), in $B_{\perp\kappa\alpha j}\{{\mathbf{R}}\}$, Eq. (19), as well as in $\hat{\cal{T}}^{\textrm{N}}_{j}\{{\mathbf{R}}\}$, Eq. (28), remain derivatives in terms of nuclear configurations. This is due to the expression of the second factor in Eq. (48), which is still expressed in nuclear configurations, not in scaled displacements. However, these functions are evaluated in the configuration expressed equivalently by $\{{\mathbf{R}}\}$ or by $\lambda\boldsymbol{\rho}]$. So, one has
@@
We now examine the mass-scaling behavior of the different approximations presented in Sec. II. Similarly to Eqs. (23a)-(23b), we express the total wavefunction using scaled displacements $\boldsymbol{\rho}_{j}]$ instead of real space coordinates:
@@
In the scalar approximation, one sets $\lambda A_{\kappa\alpha j}[\lambda\boldsymbol{\rho}_{j}]$ and $E^{\rm Born}_{j}[\lambda\boldsymbol{\rho}_{j}]$ to zero, while keeping either of them leads to alternative approximations, see Sec. LABEL:sec:hybrid of SI. The exact equation for the energy $\mathcal{E}_{i\lambda}$ of the full Hamiltonian, Eq. (58), is now expressed in terms of the self-energy, Eq. (31):
@@
In Eq. (62), the reference BO energy $E^{\rm BO}_{j|\textbf{0}}$ can be subtracted from both $E^{\rm opt}_{j\lambda}[\boldsymbol{\rho}_{j}]$ and $\mathcal{E}_{i\lambda}$, leaving quantities that are at least second-order in $\lambda$. Such second-order component can be used as a zeroth-order Hamiltonian to find the remaining expansion for $\mathcal{E}_{i\lambda}$ and to find the wavefunction $\chi_{\jmath j\lambda}[\boldsymbol{\rho}_{j}]$.
Explicitly, from Eq. (62), one deduces
$\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]\chi_{\jmath j\lambda}[\boldsymbol{\rho}_{j}]=\mathfrak{E}_{\jmath j\lambda}\chi_{\jmath j\lambda}[\boldsymbol{\rho}_{j}],$ (67)
@@
$\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]=\hat{\mathfrak{H}}^{\textrm{ho}}_{j}[\boldsymbol{\rho}_{j}]+\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}],$ (68)
that includes the harmonic oscillator Hamiltonian
$\hat{\mathfrak{H}}^{\textrm{ho}}_{j}[\boldsymbol{\rho}_{j}]\equiv\sum_{\kappa\alpha}\frac{\hat{\pi}^{2}_{\kappa\alpha}}{2m_{\kappa}}+E^{\rm BO\{\mathbf{2}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{2},$ (69)
and corrections to this harmonic oscillator Hamiltonian
$\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]\equiv$ $\lambda\hat{T}_{j}^{\textrm{vec(1)}}[\lambda\boldsymbol{\rho}_{j}]+\lambda^{2}\hat{T}_{j}^{\textrm{vec(2)}}[\lambda\boldsymbol{\rho}_{j}]$
@@
Without the corrections $\Delta\hat{\mathfrak{H}}_{j\lambda}$, Eq. (67) reduces to the $\lambda$-independent equation
$\hat{\mathfrak{H}}^{\textrm{ho}}_{j}[\boldsymbol{\rho}_{j}]\chi_{\jmath j}^{\textrm{ho}}[\boldsymbol{\rho}_{j}]=\mathfrak{E}_{\jmath j}^{\textrm{ho}}\chi_{\jmath j}^{\textrm{ho}}[\boldsymbol{\rho}_{j}],$ (75)
@@
$\chi_{\jmath j}^{\textrm{ho}}[\boldsymbol{\rho}_{j}]=\vartheta_{\jmath j}[\boldsymbol{\rho}_{j}].$ (76)
Contributions with different powers of $\lambda$ in $\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]$ can be ordered to understand the effect of ${\cal{O}}(\lambda)$ and ${\cal{O}}(\lambda^{2})$ contributions on $\mathfrak{E}_{\jmath j\lambda}$. Moreover, it is possible to fix the gauge of the BO wavefunction at the reference configuration, as well as its first-order derivative with respect to atomic displacements, such that $A_{\kappa\alpha j}[\boldsymbol{0}]=0$. Hence $\hat{T}_{j}^{\textrm{vec(1)}}$ is ${\cal{O}}(\lambda)$ and $\hat{T}_{j}^{\textrm{vec(2)}}$ is ${\cal{O}}(\lambda^{2})$. The corrections to the harmonic oscillator Hamiltonian, Eq. (III) becomes
$\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]\equiv\lambda E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}$
@@
As previously observed, $\vartheta_{\jmath j}$ is even or odd with respect to a change of sign of $\boldsymbol{\rho}_{j}$ (also changing sign of the associated momentum). Moreover, all terms of $\Delta\hat{\mathfrak{H}}_{j\lambda}$ that are even with respect to such a sign change have an even $\lambda$ exponent, while all terms that are odd with respect to such a sign change have an odd $\lambda$ exponent. Due to these symmetry considerations, the $\lambda$-expansion of the total energy turns out to have only even power terms.
The eigenenergy present in Eq. (67) is studied in the SI, Sec. LABEL:app:auxH4, and gives
$\mathcal{E}_{i}^{\rm(4)}=\mathcal{E}_{i}^{\rm(4),anh}+\mathcal{E}_{i}^{(4),\text{vec}}+E^{\rm Born}_{j}[\textbf{0}],$ (78)
with three components, first the anharmonic term,
@@
$\mathcal{E}_{i}^{\rm(4),anh}=\langle\vartheta_{\jmath j}|E^{\rm BO\{\mathbf{4}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{4}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}\\ \!+\!\langle\vartheta_{\jmath j}|\big(E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}\big)\hat{\mathfrak{G}}_{\perp\jmath j}^{\textrm{ho}}[\boldsymbol{\rho}_{j}]\big(E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}\big)|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}},$ (79)
that is present in the expansion of the scalar approximation, then also two terms that disappear in this approximation, the vector one,
$\mathcal{E}_{i}^{(4),\text{vec}}=\langle\vartheta_{\jmath j}|\hat{T}_{j}^{\textrm{vec(1)\{{\bf{1}}\}}}\cdot\boldsymbol{\rho}_{j}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}},$ (80)
and the Born one,
$E^{\rm Born}_{j}[\textbf{0}]=\langle\vartheta_{\jmath j}|E^{\rm Born}_{j}[\textbf{0}]|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}.$ (81)
In Eq. (79), we have used the Green’s function of the auxiliary harmonic oscillator Hamiltonian $\hat{\mathfrak{H}}_{j}^{\textrm{ho}}[\boldsymbol{\rho}_{j}]$, see Eq. (46). More precisely, we denote
$\hat{\mathfrak{G}}_{\perp\jmath j}(z,[\boldsymbol{\rho}_{j}])\equiv\hat{\cal{P}}_{\perp\vartheta_{\jmath j}}\Big(z-\hat{\mathfrak{H}}^{\textrm{ho}}_{j}[\boldsymbol{\rho}_{j}]\Big)^{-1}\hat{\cal{P}}_{\perp\vartheta_{\jmath j}},$ (82)
with the term between parenthesis being a pseudo-inverse, limited to the subspace perpendicular to $\vartheta_{\jmath j}[\boldsymbol{\rho}_{j}]$. The Green’s function evaluated at $z=\mathfrak{E}_{\jmath j}^{\textrm{ho}}$ is labeled with a ‘ho’ superscript, and used in Eq. (79),
$\hat{\mathfrak{G}}_{\perp\jmath j}^{\textrm{ho}}[\boldsymbol{\rho}_{j}]\equiv\hat{\mathfrak{G}}_{\perp\jmath j}(\mathfrak{E}_{\jmath j}^{\textrm{ho}},[\boldsymbol{\rho}_{j}]).$ (83)
@@
Now that we have established the mass scaling of the optimal and scalar approximations up to order five, we can carry on our search for the electron-phonon contribution to the total energy, although there is an intermediate question first: what is the phonon contribution to the total energy?